OpenStudy (anonymous):

Find the Distance between the points: (-6,5) and (2,-3) When plugged into the formula for distance, you get the square root of 0...

OpenStudy (anonymous):

\[d=\sqrt{(2+6)^2-(-3-5)^2}\]

OpenStudy (anonymous):

No, wait, I did something stupid. Nevermind.

OpenStudy (anonymous):

No, I didn't do anything stupid, i was right...

OpenStudy (anonymous):

Ugh, yes, I did something stupid...

OpenStudy (anonymous):

That subtraction should be addition.

OpenStudy (anonymous):

Yes, that, and I forgot that squaring a negative term results in a positive answer.

OpenStudy (anonymous):

This should be: \[6\sqrt{2}\]

OpenStudy (anonymous):

I got 8sqrt(2) because both those parenthesis squared are 65

OpenStudy (anonymous):

64 I mean

OpenStudy (anonymous):

8*8=64

OpenStudy (anonymous):

I broke 128 down into factors. That left me with: \[2*2*2\sqrt{6}\] You are right.

OpenStudy (anonymous):

I added them instead of multiplying them... Heh heh, ty for helping me find it though.

OpenStudy (anonymous):

No problem it would be also easy to break it down as sqrt(2*64)

OpenStudy (anonymous):

Okay yes, I follow you there.